Circular No. 6582

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 A. Guarnieri, C. Bartolini, A. Piccioni and N. Masetti, University of

Bologna; E. Costa, IAS, Frascati; F. Frontera, ITESRE and University of Ferrara; and E. Palazzi, ITESRE, Bologna, write: “A preliminary analysis of CCD observations of the GRB 970228 field (IAUC 6572), obtained with the Bologna University 1.5-m telescope in B and R bands (limiting mag ∼ 22 for both) on Feb. 28.76, Mar. 1.81, 3.76, 4.82, and 5.86 UT, shows no object fading by > 0.2 mag. A comparison between these frames and the Digital Sky Survey (DSS) shows that all of the objects brighter than B ∼ 19 and R ∼ 20 in our frames are also present in the DSS field and that their luminosities have not noticeably changed since then (1950.294).”

J. S. Bloom, N. Tanvir, and R. A. M. J. Wijers, Institute of Astronomy, Cambridge, report: “We have obtained BVR images of the error box of GRB 970228 using the 2.5-m Isaac Newton Telescope (La Palma) on Mar. 8 and 9. On Mar. 9.89 UT (seeing < 1′′), we resolve the galaxy coincident with the radio-source position (Frail et al., IAUC 6576) just outside the GRB error box (Hurley et al., IAUC 6578). We find V = 21.4, R = 21.2, B = 24.1. In the 0.7-arcmin^{2} region of overlap between the GRB error box and the error region of the MECS/LECS x-ray source (Costa et al., IAUC 6578), we find ∼ 30 objects down to B = 24.5 or R = 24. Three of the fifteen objects with R < 22 are galaxies (based on their morphology) whose α and δ end figures (equinox 2000.0), along with magnitudes (given as R, V − R, and B − V), are as follows: 46^{s}.14, 46′45′′.9, 21.6, +0.7, +0.9; 46^{s}.68, 46′17′′.0, 20.0, +1.3, +1.1; 45^{s}.41, 45′49′′.9, 20.0, +0.9, +1.7. We suspect that an object at 46^{s}.52, 47′25′′.7, 20.6, +1.4, +1.2 may also be a galaxy.”

M. R. Metzger, S. R. Kulkarni, S. G. Djorgovski, R. Gal, and C. C. Steidel, California Institute of Technology; and D. A. Frail, National Radio Astronomy Observatory, communicate: “Images of the optical counterpart (Margon et al., IAUC 6577) to the radio source detected by Frail et al. (IAUC 6576) were obtained with the Keck II 10-m telescope on Mar. 6.32 UT and with the Palomar 1.5-m reflector on Mar. 6.19, revealing an extended source of R = 20.8 with center at α = 5^{h}01^{m}41^{s}.67, δ = +11^{o}47′04′′.8 ± 0′′.2 (equinox 2000.0). This source lies in the error circle of the x-ray source seen by Costa et al. (IAUC 6576) and near the error box of GRB 970228 (IAUC 6572, 6578). A spectrum taken with Keck II on Mar. 9.25 shows a galaxy at redshift 0.498, inferred from Ca II absorption and narrow [O II] 372.8-nm and [O III] 500.7-nm emission features.”

1997 March 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green